Hi

What are the products of the following reaction:

B+LiNO3+C + H2O ——->

Thanks
1
Expert's answer
2018-10-04T07:14:41-0400
2B + 3H2O = B2O3 + 3H2
C + H2O = CO + H2
LiNO3 + 4H2O = [Li(H2O)4]+ + NO3-
LiNO3 + 2B + C + 8H2O = [Li(H2O)4]NO3 + B2O3 + CO + 4H2
t = 700-9000C